OpenStudy (aaronandyson):

The diagonal of a rectangle is 60 m more than its shorter side and the larger side is 30 m more than its shorter side.Find the sides of the rectangle @hartnn

hartnn (hartnn):

so let shorter side be 'x' larger side = x+30 diagonal = x+60 then pythagoras theorem

OpenStudy (aaronandyson):

Is (x+60)^2 the same as (60+x)^2?

hartnn (hartnn):

yes :)

hartnn (hartnn):

\(x^2 + (x+30)^2 = (x+60)^2\) get x

OpenStudy (aaronandyson):

x^2 + (x^2 + 900 + 60x) = (x^2 + 3600 + 120) ?

hartnn (hartnn):

120x and yes

OpenStudy (aaronandyson):

The numbers are killing me X_X

hartnn (hartnn):

also side length cannot be negative

OpenStudy (aaronandyson):

I did it without a calci I got x^2 - 60x - 2700 = 0?

OpenStudy (aaronandyson):

I'm right?

hartnn (hartnn):

yes

OpenStudy (aaronandyson):

Again, Help me breaking the terms now....

hartnn (hartnn):

27 = 3*9 2700 = 30*90

hartnn (hartnn):

-60x = -90x + 30x

OpenStudy (aaronandyson):

x^2 - 90x + 30x - 2700 = 0?

hartnn (hartnn):

yep, find 2 values of x , then eliminate one

OpenStudy (aaronandyson):

Is it 90 cm ?

hartnn (hartnn):

yes! shorter = 90 larger = 90+30 :)

OpenStudy (aaronandyson):

120 cm !

hartnn (hartnn):

not cm 90m and 120m :)

OpenStudy (aaronandyson):

Typos X_X

hartnn (hartnn):

maybe the question had typos, and it really is cm, and you're correct :P

OpenStudy (aaronandyson):

I need to do 8 more problems ....

hartnn (hartnn):

good luck, have fun!

OpenStudy (aaronandyson):

Can you help me again???